A sample of copper contains two isotopes of relative isotopic mass 63.0 and 65.0. If the relative
atomic mass of copper is 63.5, calculate the relative abundance of each isotope.
1
Expert's answer
2020-10-20T08:36:14-0400

Solution:

Let X be the relative abundance of the isotope with relative mass 63.0. Then the relative abundance of the 65.0 isotope is (1-X). Thus:

"63.0\times{X}+65.0\times(1-X)=63.5"

"65.0-2\times{X}=63.5"

"X=0.75"

The relative abundance of the isotope with mass 63.0 is "0.75=75\%" . Therefore, the relative abundance of the 65.0 isotope is "(1-0.75)=0.25=25\%" .


Answer:

The relative abundance of the isotope with mass 63.0 is "0.75=75\%" . Therefore, the relative abundance of the 65.0 isotope is "0.25=25\%" .